OpenStudy (anonymous):

I don't understand this PLEASE HELP! I will give a medal!!! Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) greater than or equal to -31? x greater than or equal to -9 -9 less than or greater to x A number line with a closed circle on -9 and shading to the left A number line with a closed circle on -9 and shading to the right

OpenStudy (anonymous):

Check to see if you have written the second answer correctly. Should it read like this: -9 less than or equal to x? \[8x-(5x+4)\ge-31\]\[8x-5x-4\ge-31\]\[3x-4\ge-31\]\[3x \ge-27\]\[x \ge-9\] Answer one is the right way Answer four is the right way Answer three is the wrong way Answer two a typo? See this web site http://www.mathsisfun.com/sets/intervals.html